OpenStudy (anonymous):

At what points are the equations y=x2 and y=1/x+2 equal?

OpenStudy (loser66):

if you want them equal, then let them equal and solve for x, then y. Done

OpenStudy (z4k4r1y4):

First equation: y = something , second equation: y = something else. So you can have: something = something else.

OpenStudy (loser66):

Hey, feather, this site is so boring, right? students just want the free answer. Wonder why no one wants to study!!

OpenStudy (anonymous):

Medals to anyone that can help!

OpenStudy (shamim):

Can we write x^2=1/(x+2)

OpenStudy (anonymous):

Yeah I made a typing error

OpenStudy (shamim):

x^2(x+2)=1 Right?

OpenStudy (anonymous):

Yes, I wrote that down.

OpenStudy (shamim):

@Losser66 help plz!!

OpenStudy (shamim):

@Loser66